# Tree-Like Shortcuttings of Trees

### v1

@@ 4 Bounded arboricity tree covers
4.1 Line metrics
In this section, we show a construction for line metrics (Section˜1.1.1). We shall use a modification of the following well-known result.
@@ 4.1 Line metrics
For every $n\geq 2$ and $k\geq 2$, every $n$-point tree metric admits a 1-spanner with hop-diameter $k$ and $O(n\alpha_{k}(n))$ edges.
@@
Let $n\geq 2$ be an arbitrary integer. Let $L$ be a line metric induced by a set of $n$ points on a line so that between every two points there is $n$ Steiner points. Let $S$ denote the set of Steiner points. Then, $L$ admits a Steiner $1$-spanner with hop-diameter $2$ such that the Steiner points belong to $S$ and every vertex in $L\cup S$ has a constant in-degree.
@@
Let $n\geq 2$ and $k\geq 2$ be two arbitrary integers. Let $L$ be a line metric induced by a set of $n$ points on a line so that between every two points there is $\alpha_{k}(n)$ Steiner points. Let $S$ denote the set of Steiner points. Then, $L$ admits a Steiner $1$-spanner with hop-diameter $2k$ such that the Steiner points belong to $S$ and every vertex in $L\cup S$ has a constant in-degree.
@@
We prove the lemma by induction over $k$. For $k=2$, we take a central vertex $c$ in $L$ and connect it to every other point in $L$; orient the edges outwards from $c$. Proceed recursively with the two halves. This way we obtain a $1$-spanner for $L$ with hop-diameter $2$. Denote by $E$ the edge set of this spanner. The depth of the recursion in the construction is $O(\log{n})$, and the size of $S$ is $n\alpha_{2}(n)=n\log{n}$. Every edge in $E$ has in-degree 1 for every recursion level. We can split every such edge into two using a fresh Steiner point for each recursion level. This concludes the proof for $k=2$.
@@
[…]
@@
Let $L_{n}$ be an arbitrary line metric. For an integer $k^{\prime}\geq 2$, we describe a construction of a 1-spanner $H$ for $L_{n}$ with hop-diameter $4k^{\prime}-2$ and arboricity $O(\alpha_{2k^{\prime}}(n))$.
Consider a set of $n^{\prime}=n/\alpha_{2k^{\prime}-2}(n)$ equally-spaced cut vertices dividing $L_{n}$ into intervals of size $\alpha_{2k^{\prime}-2}(n)$. To construct the 1-spanner $H$, we connect every cut vertex to all the vertices in its two neighboring intervals. Denote the corresponding edge by $E_{C}$. Let $C$ be the set of cut vertices. Let $E^{\prime}$ be obtained by invoking Lemma˜4.3 with parameter $2k^{\prime}-2$ on $C$ using $L_{n}$ as Steiner points. Proceed recursively with each of the intervals.
To analyze the arboricity, we will show that the edges in $E_{C}$ and $E^{\prime}$ can be oriented so that the in-degree of every vertex is constant. Orient every edge in $E_{C}$ so that it goes out of the corresponding cut vertex. Since every interval is adjacent to at most two cut vertices, the in-degree of every point with respect to $E_{C}$ is at most 2. By Lemma˜4.3, the edges in $E^{\prime}$ have a constant in-degree on $L_{n}$. In conclusion, $E_{C}$ and $E^{\prime}$ contribute $O(1)$ to the in-degree of each vertex in $L_{n}$. The number of recursion levels $\ell(n)$ satisfies the recurrence $\ell(n)=\ell(\alpha_{2k^{\prime}-2}(n))+O(1)$, which has solution $\ell(n)=\alpha_{2k^{\prime}}(n)$. Every recursion level contributes $O(1)$ to the in-degree of vertices, meaning that the overall in-degree in $H$ is $O(\alpha_{2k^{\prime}}(n))$. The hop-diameter of $H$ is $2+2\cdot(2k^{\prime}-2)=4k^{\prime}-2$. This concludes the description of a 1-spanner with hop-diameter $4k^{\prime}-2$ and arboricity $O(\alpha_{2k^{\prime}}(n)$.
@@
The construction for ultrametric is a simple adaptation of the construction for line metrics, which we will show next.
@@ 4.2 Ultrametrics and doubling metrics
Chan and Gupta [CG06] showed that any $(1/\epsilon,2)$-HST can be embedded into the line metric with (worst-case) distortion $1+O(\epsilon)$. Therefore, by applying Section˜1.1.1, we obtain a $(1+O(\epsilon))$-spanner with hop diameter $k$ and arboricity $O(\alpha_{k/2+1}(n))$ for $(1/\epsilon,2)$-HST. In our setting, we are interested in large-degree $(k,\Delta)$-HST where $s=1/\epsilon$ and $\Delta=\operatorname{\mathrm{poly}}(1/\epsilon)$; the embedding result by Chan and Gupta [CG06] no longer holds for these HSTs. Instead, we directly apply our technique for the line metric to get a $(1+O(\epsilon))$-spanner with low-hop diameter.
@@
Let $T$ be the $(1/\epsilon,\Delta)$-HST and let $M_{T}$ be the metric induced by $T$. For an integer $k^{\prime}\geq 2$, we describe a construction of a $(1+O(\epsilon))$-spanner for $M_{T}$ with hop-diameter $4k^{\prime}-2$ and arboricity $O(\alpha_{2k^{\prime}}(n))$. The construction is similar to that in Section˜1.1.1.
@@
The same argument in Section˜1.1.1 implies that the arboricity is $O(\alpha_{2k^{\prime}}(n))$. The stretch is $(1+O(\epsilon))$ since the path $u\rightarrow x\rightarrow v$ between two cut vertices $u$ and $v$ has stretch $(1+O(\epsilon))$. ∎
To construct a low-hop spanner with small arboricity for doubling metrics (Section˜1.1.2), we will rely on the ultrametric cover by Filtser and Le [FL22a]. Following their notation, for a given metric space $(X,d_{X})$, we say that a collection $\mathcal{T}$ of at most $\tau$ different $(s,\Delta)$-HSTs such that (i) for every HST $T\in\mathcal{T}$, points in $X$ are leaves in $T$, and (ii) for every two points $x,y\in X$, $d_{X}(x,y)\leq d_{T}(x,y)$ for every $T\in\mathcal{T}$, and there exists a tree $T_{xy}\in\mathcal{T}$ such that $d_{X}(x,y)\leq\rho\cdot d_{T_{xy}}(x,y)$.
@@ 4.3 General tree metrics
[…]
Proof of Section˜1.1.1 for height $h$.
@@
Proof of Section˜1.1.1 for general trees.
@@
We start by employing the shortcutting procedure for bounded height trees on $T^{\prime}$ with parameter $k^{\prime}$ and explain how to adapt it to $T$. Consider an arbitrary edge $(u,v)$ in the shortcutting of $T$ such that $u$ is a descendant of $v$. Let $P_{u}$ (resp., $P_{v}$) be the heavy path in $T$ corresponding to $u$ (resp., $v$). Add an edge between the root $r_{u}$ of $P_{u}$ and its lowest ancestor on $P_{v}$. We do this for all the edges in the shortcutting of $T^{\prime}$. For every heavy path $P$ in $T$, use the 4-hop construction with arboricity $O(\log\log{n})$ from Section˜1.1.1. In addition, connect via a direct edge every vertex in $p$ to the root of $P$. This concludes the description of the shortcutting for $T$.
@@
5 Routing schemes
@@ 5.1 Routing in tree metrics
[…]
@@ 5.2 Routing in doubling metrics
We equip each tree in the cover with the stretch-1 routing scheme from Theorem˜1.2. This consumes overall $\epsilon^{-\tilde{O}(d)}\log^{2}n/\log\log{n}$ bits per point in $P$. In addition, we add $\epsilon^{-\tilde{O}(d)}\log{n}$-bit labels to each point in $P$ as stated in Lemma˜5.1. Given two points $x,y\in P$, we first employ the algorithm from Lemma˜5.1 to find the tree in which the routing should proceed. Then, the routing proceeds on that specific tree using the routing algorithm from Theorem˜1.2. This concludes the description of the compact routing scheme for doubling metrics.
6 Lower bound for routing in tree metrics
[…]

### v2

@@ 4 Low arboricity shortcuttings
4.1 Paths
In this section, we show a construction for paths (Lemma˜4.3). We shall use a modification of the following well-known result.
@@ 4.1 Paths
For every $n\geq 2$ and $k\geq 2$, every $n$-point tree admits a shortcutting with hop-diameter $k$ and $O(n\alpha_{k}(n))$ edges.
@@
Let $n\geq 2$ be an arbitrary integer. Let $L$ be a path induced by a set of $n$ points on a line so that between every two points there is $n$ Steiner points. Let $S$ denote the set of Steiner points. Then, $L$ admits a Steiner shortcutting with hop-diameter $2$ such that the Steiner points belong to $S$ and every vertex in $L\cup S$ has a constant in-degree.
@@
Let $n\geq 2$ and $k\geq 2$ be two arbitrary integers. Let $L$ be a path induced by a set of $n$ points on a line so that between every two points there is $\alpha_{k}(n)$ Steiner points. Let $S$ denote the set of Steiner points. Then, $L$ admits a Steiner shortcutting with hop-diameter $2k$ such that the Steiner points belong to $S$ and every vertex in $L\cup S$ has a constant in-degree.
@@
We prove the lemma by induction over $k$. For $k=2$, we take a central vertex $c$ in $L$ and connect it to every other point in $L$; orient the edges outwards from $c$. Proceed recursively with the two halves. This way we obtain a shortcutting for $L$ with hop-diameter $2$. Denote by $E$ the edge set of this shortcutting. The depth of the recursion in the construction is $O(\log{n})$, and the size of $S$ is $n\alpha_{2}(n)=n\log{n}$. Every edge in $E$ has in-degree 1 for every recursion level. We can split every such edge into two using a fresh Steiner point for each recursion level. This concludes the proof for $k=2$.
@@
We next prove Lemma˜4.3.
{restatable}
theoremarbUBline For every $n\geq 1$ and every even $k\geq 2$, every $n$-point path admits a shortcutting with hop-diameter $k$ and arboricity $O(\alpha_{k/2+1}(n))$.
Figure 3: An illustration of the shortcutting with hop-diameter 4 and arboricity $O(\log\log{n})$. We start by a well-known 3-hop construction [AS87, BTS94]. Split the path into subpaths using $\sqrt{n}$ cut vertices (red) so that every subpath contains $\sqrt{n}$ vertices. The shortcutting is obtained by: (i) interconnecting the cut vertices by a clique, (ii) connecting every cut vertex to the points in its two adjacent subpaths (in the illustration, only two such edges, $(u^{\prime},u)$ and $(v,v^{\prime})$, are shown in green), and (iii) applying the recursive construction to each of the subpaths. This construction readily gives hop-diameter 3 (see, e.g., the path between $u^{\prime}$ and $v^{\prime}$), but it suffers from having arboricity $\Omega(\sqrt{n})$ due to the clique on the cut vertices. To remedy this, we split every clique edge $(u,v)$ into two edges, $(u,w)$ and $(w,v)$. This allows us to reduce the in-degree of every point to $O(1)$ per recursion level, at a cost of increasing the hop-diameter to 4. Overall, the number of recursion levels is $O(\log\log{n})$, yielding the desired bound on the arboricity.
@@
Let $L_{n}$ be an arbitrary path. For an integer $k^{\prime}\geq 2$, we describe a construction of a shortcutting $H$ for $L_{n}$ with hop-diameter $4k^{\prime}-2$ and arboricity $O(\alpha_{2k^{\prime}}(n))$.
Consider a set of $n^{\prime}=n/\alpha_{2k^{\prime}-2}(n)$ equally-spaced cut vertices dividing $L_{n}$ into intervals of size $\alpha_{2k^{\prime}-2}(n)$. To construct the shortcutting $H$, we connect every cut vertex to all the vertices in its two neighboring intervals. Denote the corresponding edge by $E_{C}$. Let $C$ be the set of cut vertices. Let $E^{\prime}$ be obtained by invoking Lemma˜4.3 with parameter $2k^{\prime}-2$ on $C$ using $L_{n}$ as Steiner points. Proceed recursively with each of the intervals.
To analyze the arboricity, we will show that the edges in $E_{C}$ and $E^{\prime}$ can be oriented so that the in-degree of every vertex is constant. Orient every edge in $E_{C}$ so that it goes out of the corresponding cut vertex. Since every interval is adjacent to at most two cut vertices, the in-degree of every point with respect to $E_{C}$ is at most 2. By Lemma˜4.3, the edges in $E^{\prime}$ have a constant in-degree on $L_{n}$. In conclusion, $E_{C}$ and $E^{\prime}$ contribute $O(1)$ to the in-degree of each vertex in $L_{n}$. The number of recursion levels $\ell(n)$ satisfies the recurrence $\ell(n)=\ell(\alpha_{2k^{\prime}-2}(n))+O(1)$, which has solution $\ell(n)=\alpha_{2k^{\prime}}(n)$. Every recursion level contributes $O(1)$ to the in-degree of vertices, meaning that the overall in-degree in $H$ is $O(\alpha_{2k^{\prime}}(n))$. The hop-diameter of $H$ is $2+2\cdot(2k^{\prime}-2)=4k^{\prime}-2$. This concludes the description of a shortcutting with hop-diameter $4k^{\prime}-2$ and arboricity $O(\alpha_{2k^{\prime}}(n)$.
@@
The construction for ultrametric is a simple adaptation of the construction for paths, which we will show next.
@@ 4.2 Ultrametrics and doubling metrics
Chan and Gupta [CG06] showed that any $(1/\epsilon,2)$-HST can be embedded into the line metric with (worst-case) distortion $1+O(\epsilon)$. Therefore, by applying Lemma˜4.3, we obtain a $(1+O(\epsilon))$-spanner with hop diameter $k$ and arboricity $O(\alpha_{k/2+1}(n))$ for $(1/\epsilon,2)$-HST. In our setting, we are interested in large-degree $(k,\Delta)$-HST where $s=1/\epsilon$ and $\Delta=\operatorname{\mathrm{poly}}(1/\epsilon)$; the embedding result by Chan and Gupta [CG06] no longer holds for these HSTs. Instead, we directly apply our technique for the line metric to get a $(1+O(\epsilon))$-spanner with low-hop diameter.
@@
Let $T$ be the $(1/\epsilon,\Delta)$-HST and let $M_{T}$ be the metric induced by $T$. For an integer $k^{\prime}\geq 2$, we describe a construction of a $(1+O(\epsilon))$-spanner for $M_{T}$ with hop-diameter $4k^{\prime}-2$ and arboricity $O(\alpha_{2k^{\prime}}(n))$. The construction is similar to that in Lemma˜4.3.
@@
The same argument in Lemma˜4.3 implies that the arboricity is $O(\alpha_{2k^{\prime}}(n))$. The stretch is $(1+O(\epsilon))$ since the path $u\rightarrow x\rightarrow v$ between two cut vertices $u$ and $v$ has stretch $(1+O(\epsilon))$. ∎
To construct a low-hop spanner with small arboricity for doubling metrics (Section˜1.1), we will rely on the ultrametric cover by Filtser and Le [FL22a]. Following their notation, for a given metric space $(X,d_{X})$, we say that a collection $\mathcal{T}$ of at most $\tau$ different $(s,\Delta)$-HSTs such that (i) for every HST $T\in\mathcal{T}$, points in $X$ are leaves in $T$, and (ii) for every two points $x,y\in X$, $d_{X}(x,y)\leq d_{T}(x,y)$ for every $T\in\mathcal{T}$, and there exists a tree $T_{xy}\in\mathcal{T}$ such that $d_{X}(x,y)\leq\rho\cdot d_{T_{xy}}(x,y)$.
@@ 4.3 General tree metrics
{restatable}
theoremarbUBtree For every two integers $n\geq 1$ and $k\geq 1$ and every $n$-vertex tree $T$, there is a shortcutting with hop-diameter $k$ and arboricity $O(\log^{12/(k+4)}{n})$. Moreover, when the height of the tree is $h$, then the arboricity is $O(h^{6/(k+4)})$.
Proof of Section˜4.3 for height $h$.
@@
Proof of Section˜4.3 for general trees.
@@
We start by employing the shortcutting procedure for bounded height trees on $T^{\prime}$ with parameter $k^{\prime}$ and explain how to adapt it to $T$. Consider an arbitrary edge $(u,v)$ in the shortcutting of $T$ such that $u$ is a descendant of $v$. Let $P_{u}$ (resp., $P_{v}$) be the heavy path in $T$ corresponding to $u$ (resp., $v$). Add an edge between the root $r_{u}$ of $P_{u}$ and its lowest ancestor on $P_{v}$. We do this for all the edges in the shortcutting of $T^{\prime}$. For every heavy path $P$ in $T$, use the 4-hop construction with arboricity $O(\log\log{n})$ from Lemma˜4.3. In addition, connect via a direct edge every vertex in $p$ to the root of $P$. This concludes the description of the shortcutting for $T$.
@@
Using Section˜4.3 and the known tree cover constructions, we obtain low arboricity spanners for various metric spaces. See Table˜1.
metric stretch arboricity
planar [CCL+23] $1+\epsilon$ $O(\epsilon^{-3}\log(1/\epsilon)\log^{12/(k+4)}{n})$
minor-free [CCL+24a] $1+\epsilon$ $2^{r^{O(r)}/\epsilon}\cdot\log^{12/(k+4)}{n}$
general [MN07] $O(\gamma)$ $O(\gamma\cdot n^{1/\gamma}\cdot\log^{12/(k+4)}{n})$
Table 1: Low-arboricity spanners for various metric spaces using Section˜4.3.
5 Routing upper bounds
@@ 5.1 Routing in tree metrics
{restatable}
theoremroutingUB For every $n$ and every $n$-vertex tree $T$, there is a 3-hop routing scheme in the fixed-port model for the metric $M_{T}$ induced by $T$ with stretch 1 that uses $O(\log^{2}{n}/\log\log{n})$ bits per vertex.
@@ 5.2 Routing in doubling metrics
We equip each tree in the cover with the stretch-1 routing scheme from Section˜5.1. This consumes overall $\epsilon^{-\tilde{O}(d)}\log^{2}n/\log\log{n}$ bits per point in $P$. In addition, we add $\epsilon^{-\tilde{O}(d)}\log{n}$-bit labels to each point in $P$ as stated in Lemma˜5.1. Given two points $x,y\in P$, we first employ the algorithm from Lemma˜5.1 to find the tree in which the routing should proceed. Then, the routing proceeds on that specific tree using the routing algorithm from Section˜5.1. This concludes the description of the compact routing scheme for doubling metrics.
6 Routing lower bound
In this section, we prove the following theorem. {restatable}theoremroutingLB There is an infinite family of trees $T_{n}$, $n>0$, such that any labeled fixed-port routing scheme with stretch 1 on a metric induced by $T_{n}$ has at least one vertex with total memory of $\Omega(\log^{2}{n}/\log\log{n})$ bits.
@@ 6 Routing lower bound
Acknowledgement.
Hung Le and Cuong Than are supported by NSF grant CCF-2517033 and NSF CAREER Award CCF-2237288. Cuong Than is also supported by a Google Ph.D. Fellowship. Shay Solomon is funded by the European Union (ERC, DynOpt, 101043159). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Lazar Milenković and Shay Solomon are funded by a grant from the United States-Israel Binational Science Foundation (BSF), Jerusalem, Israel, and the United States National Science Foundation (NSF).